OpenStudy (anonymous):

In the products 2axla =4 1a ; 2bx1b=37b, a and b represent positive integers. Then a+b=

OpenStudy (aravindg):

What does axla represent?

OpenStudy (homeworksucks):

Is this advanced math that I haven't covered yet, or is the l not supposed to be there

OpenStudy (anonymous):

pls send the answer

OpenStudy (aravindg):

Yes. Sure but please clarify to my question above.

OpenStudy (anonymous):

this is advanced maths?

OpenStudy (aravindg):

What does axla represent?

OpenStudy (anonymous):

i can't understand. what is the meaning of axla?

OpenStudy (aravindg):

2axla =4 1a What does this represent? I am not getting your notation.

OpenStudy (anonymous):

2a x 1b = 41 a

OpenStudy (homeworksucks):

\[2a \times 1b=41a\]

OpenStudy (aravindg):

That makes sense now.

OpenStudy (aravindg):

So we are dealing with vectors isnt it?

OpenStudy (anonymous):

In the products 2a x la =41a ; 2b x 1b=37b, a and b represent positive integers. Then a+b= ?

OpenStudy (anonymous):

This question is presented in the government exam questions in group 4

OpenStudy (aravindg):

I will assume these are scalars 2a x a=41a 2a=41 a=?

OpenStudy (anonymous):

k then

OpenStudy (aravindg):

Similary 2b=37 Solve for b Then just do a+b

OpenStudy (aravindg):

*similarly

OpenStudy (anonymous):

pls send the answer a+b?

OpenStudy (aravindg):

We just solved for a and b above. Just add them!

OpenStudy (aravindg):

Got it?

OpenStudy (anonymous):

sry. u reply a+b?

OpenStudy (aravindg):

Sorry in this website we do not give you direct answers instead just help you et to the answer. Please do the last step its easier than you can think!

OpenStudy (aravindg):

*get to the answer

OpenStudy (aravindg):

What did you get from this: 2a x a=41a 2a=41 a=?

OpenStudy (anonymous):

20.5

OpenStudy (aravindg):

Great! Similarly what did you get for b?

OpenStudy (anonymous):

18.5

OpenStudy (aravindg):

Yes! Now simply add them?

OpenStudy (anonymous):

k now this is the answer is 39. But in the answer key is 11. how is it?

OpenStudy (aravindg):

11?? Are you sure you typed the question correctly?

OpenStudy (anonymous):

yes. In the products 2a x la =41a ; 2b x 1b=37b, a and b represent positive integers. Then a+b= ? this is the question.

OpenStudy (aravindg):

It should be 39.

OpenStudy (aravindg):

Maybe the key is wrong?

OpenStudy (anonymous):

k

OpenStudy (aravindg):

Okay So good luck!

OpenStudy (anonymous):

In an examination 30% of total students failed in English, 40% of students failed in Hindi and 20% in bot'h. Find the percentage of students who passed in both the subjects.

OpenStudy (aravindg):

Please post your questions separately. Use the close button above and post a new question.

OpenStudy (anonymous):

In the products 2axla =4 1a ; 2bx1b=37b, a and b represent positive integers. Then a+b= here positive integers are those after 0 ie 1,2,3,4,5,6,7,8,9 so substitute these numbers in the place of a and b so 2a*1a=41a here substitute 6 ie 26*16=416 so a=6 then 2b*1b=37b here substitute 5 ie 25*15=375 so b=5 then a+b=11 ie 6+5=11.